Supplemental Operations Report 31-DEC 2022

Supplemental Operations Report 31-DEC-2022

Name of person filing report: Shannon Rupert

Reason for Report: Routine/ Last day of 2022!

Non-nominal systems: Power system, toilet, cabinet door under sink in Hab

Action taken for non-nominal systems: See notes below for toilet and power system. The cabinet door might not be replaceable and if it isn’t we will go forage through the storage and see if there is an extra one in the cabinets for the ScienceDome

Generator: Running from about 8 pm to 8 am. The generator was repaired on Thursday. The problem was the governor, which had lost a small plastic pin and as a result was unstable and the vibration of the engine was causing the surging. How crazy! We also needed the oil pressure sensor changed as it was faulty and we had the bushings replaced because they would have needed it soon. The technician was great and spent a lot of time showing us what things were that would commonly need replaced and how to replace them. I am very grateful to not have to worry every night about losing power! The oil pressure sensor will be mailed to us, but it is easy to replace. He did not have one with him.

Solar— Nominal but we have had some cloudy days and since the generator currently won’t charge the batteries, we have had to use the generator more hours than normal.

Solar— SOC Last 24 hours:

Max 100

Min 44

Avg 65.9

Power system notes: I reprogramed the router so SOC can again be monitored remotely. Yay, me!

Propane Reading, station tank – 73 % (all propane tanks were filled on Tuesday, the 27th, in anticipation of the storm

Propane Reading, director tank— 83 %

Propane Reading, intern tank— 81 %

Propane Reading, generator— 73 %

Ethanol Free Gasoline – 0 gallons

Water (static tank) – 550 gallons

Water in GreenHab – not noted gallons

Water in ScienceDome: 0 gallons

Water (Outpost tank) – 400 gallons

Hab toilet tank emptied: See crew operations reports. On the 31st, I did flush the tank three times: 1. Filled it up and emptied what was left from the crew. It smelled. 2. Added a bottle of dishsoap and 5 gallons of hot water and emptied. 3. Filled it with fresh water and emptied, then added water and digester so it is usable. Each time I emptied it I still heard the glub, glub, meaning something is still wrong (how do you think my very professional words?)

Sojourner rover used: yes

Hours: not noted

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Notes on rovers: nothing to report

ATV’s Used: (Honda, 350.1, 350.2, 300): none

Reason for use: n/a

Oil Added? no

ATV Fuel Used: 0 Gals

# Hours the ATVs were Used today: 0

Notes on ATVs: Nothing to report

HabCar used and why, where? Yes, to town

CrewCar used and why, where? Yes, to town and to Grand Junction

Luna used and why, where? No

General notes and comments: Nothing to note

Summary of internet: Nominal. HughesNet is being moved from my trailer and will be reinstalled in the Hab for crew use during simulation.

Summary of suits and radios: See crew operations reports.

EVA COMMS: Further changes to the COMMS system did not improve signal. They were: all radios with larger antenna, a rover mounted unit, and extending the repeater antenna by an additional eight feet. The decision was made to reinstall the system on the North Ridge, which was where it was when we tested it last season. Crew 271 did find a route to the top of the ridge, but it is going to take a huge effort to move it. We will do this as soon as we can, but it may be awhile before everything comes together.

Campus wide inspection, if action taken, what and why? RAM exterior needs some repainting.

Summary of general operations: It was so strange to have two external people on site in one week to work on systems. The crew mentioned this and I hadn’t thought about it, but I think I’m so lucky to have found people who can help when we really need it. We had the generator technician and then the electrician and the off-grid guy.

Summary of Hab operations: One of the bean bag chairs broke and the crew threw it away although I had planned to send it back. One of the two doors on the kitchen cabinet under the sink boke and is scheduled to be repaired tomorrow. We noted several things needing attention during the two week period at the end of January when we have no crew. The electrician was also briefed on where and what we wanted done with outlets, etc. He will determine where electrical rewiring needs to be done. The off-grid guy studied the power system and we shall see if he can solve the mystery of why the generator and the inverters won’t get along.

Summary of Outpost operations: General cleanup in preparation for upgrading deck.

Summary of GreenHab operations: Supplemental light 10-2 pm.

Summary of ScienceDome operations: Standing desk was returned from the Hab.

ScienceDome Dual Split: on at 65 degrees overnight (when I remember).

Summary of RAM operations: We have a lot of used oil to recycle.

Summary of any observatory issues: Both observatories were used by Crew 271 and some issues were reported to Peter.

Summary of health and safety issues: Nothing to report

Questions, concerns, supplies needed and requests: Happy New Year and welcome to Donald Jacques, our new Site Manager, who arrived at MDRS this evening in his big black bus!

Copyright © The Mars Society. All rights reserved. | Main Site